# Distributed Inter-Strand Coupling Current Model for Finite Element Simulations of Rutherford Cables

### v1

@@ 3.2 Nonlinear model with strand internal dynamics
The relation between strand current and voltage consists of two parts. The first part accounts for nonlinear inductive effects and is equal to the time derivative of the hysteretic internal magnetic flux $\Phi$ stored in the strand. The relationship between the strand current $I_{i}$ and $\Phi$ is described by the ROHF model [27]. It is similar to the ROHM model, but it defines a scalar relationship and does not contain a branch for IF coupling currents, as these can be neglected for pure transport current excitations. The ROHF model contributes to the hysteresis and eddy loss. We refer to B.2 and [27] for details about the ROHF model and its implementation.
The second part of the current to voltage relationship reproduces resistive effects due to current sharing [33] between saturating superconducting filaments, whose resistivity is described by the power law [44], and the normal conducting strand. The current sharing solution [45] defines a nonlinear resistive contribution $E(I_{i})$ to the strand voltage, for a given current $I_{i}$. It is associated with the ohmic loss contribution. See B.3 for an example.
Introducing the ROHM and ROHF models, as well as the current sharing model for the transport current, the formulation in Eq. (3.1) can be generalized as
@@
This problem is nonlinear and cannot be solved in the frequency domain anymore. An implicit Euler scheme is used for time-stepping, and a Newton-Raphson technique is implemented as an iterative scheme. The Jacobian tensor for the ROHM model is derived in [16] and the derivatives $\text{d}\Phi/\text{d}I$ and $\text{d}E/\text{d}I$ can be easily derived analytically. The convergence criterion for the Newton-Raphson scheme is a condition on the relative change of the instantaneous power computed for two successive iterations.
@@ 4 Homogenized cable model
In this section, we introduce the homogenized cable model. In this model, the cable is no longer modelled with individual strands and circuit connections for IS coupling currents, as in Section 3, but rather as a homogeneous material described with differential equations that reproduce the IS coupling current dynamics, see Figs. 1(f) and 6. The effect of contacts between strands is modelled in a distributed manner, with the Distributed Inter-Strand Coupling Current (DISCC) model. The key assumption of the homogenized model is that strands are numerous enough for a description in terms of distributed field quantities to be sufficiently accurate.
Figure 6: FE meshes for reference and homogenized cable models. The homogenized cable domain $\Omega_{\text{c}}$ does not include the artificial coating domain $\Omega_{\text{coat}}$ of the reference model. The homogenized cable domain is the smallest rectangle (or trapezoid, see A) that contains all the bare strands, and the gaps between them, as represented in red in both subfigures.
For the homogenized cable model, equations are written in terms of homogenized fields, which are spatially averaged versions of the fields in the reference cable model. For simplicity, however, we keep the same notations for the fields as before, e.g., in this section, $\boldsymbol{h}$ and $\boldsymbol{j}$ refer to the homogenized magnetic field and current density.
Because contacts between strands are treated in a distributed way in the homogenized cable model, we introduce contact resistivities $r_{\text{a}}$ and $r_{\text{c}}$ ($\Omega\,$m2), defined as the contact resistances multiplied by characteristic surfaces:
@@
We first describe the DISCC model in a linear setting in Section 4.1, focussing only on the IS coupling currents. We then generalize the model to a nonlinear setting in Section 4.2 by including hysteresis, eddy, and IF coupling current effects via distributed versions of the ROHM, ROHF, and current sharing models. The accuracy and computational performance of the linear and nonlinear models are discussed in detail in Sections 5 and 6, respectively.
@@ 4.1 Linear model for distributed IS coupling currents
We start in a linear setting in which strands are considered to be normal-conducting and non-magnetic materials with negligible resistivity for axial currents. The IS coupling currents constitute the only loss contribution in this case. The other loss mechanisms are introduced in Section 4.2.
@@ 4.1.1 Strong form of the DISCC model
For simplicity, we consider flat cables with a rectangular cross-section. We also assume uniform adjacent $r_{\text{a}}$ and crossing $r_{\text{c}}$ contact resistivities in the cable. These simplifications are however not necessary for the model, we discuss how to handle keystone angle effects [46] and non-uniform resistivities in A.
In absence of IS coupling currents, every strand carries the same net transport current (we assume that the current is uniformly shared among the strands at the termination joints), such that the homogenized current density $\boldsymbol{j}$ must therefore be uniform.
@@
Figure 7: Two excitation modes for the IS coupling currents, with external field ramp applied along (a) $\hat{\boldsymbol{e}}_{y}$, or (b) $\hat{\boldsymbol{e}}_{x}$. This illustrates that, in first approximation, (a) a variation of the strand currents along $\hat{\boldsymbol{e}}_{x}$ is associated with crossing IS coupling currents along $\hat{\boldsymbol{e}}_{y}$, and that (b) a difference of the strand currents between both layers (variation along $\hat{\boldsymbol{e}}_{y}$) is associated with adjacent IS coupling currents along $\hat{\boldsymbol{e}}_{x}$ (strands in the top schematics are represented without adjacent contacts for better visualization). This justifies the use of the curl operator on the axial current density to describe IS coupling currents.
It is important to mention that Eq. (7) is only an approximation. Indeed, non-zero adjacent IS coupling currents (along $\hat{\boldsymbol{e}}_{x}$) also take place for magnetic field applied along $\hat{\boldsymbol{e}}_{y}$, as can be seen in Fig. 5 and as discussed in [47, 26]. We however keep Eq. (7) as an assumption, and we account for these adjacent IS currents later in this section (via the mixing parameter $\lambda_{\text{a}}$ in Eq. (15)). The accuracy of the results is discussed in Section 5.
Moreover, Eq. (7) has to be seen as a conceptual relationship, which encodes the geometrical periodicity of the cable in a distributed manner. When electromagnetic fields are treated as differential forms [48, 49], it becomes evident that the curl is not the proper operator to apply on the current density (2-form). For now, we let aside these considerations and we focus instead on presenting intermediate steps leading to the weak form of the problem in Section 4.1.2. In the final weak form, only valid operators are applied on differential forms, i.e., Eq. (7) does not appear explicitly.
To close the mathematical problem, we must now express, in a distributed way, how the IS coupling current density field $\boldsymbol{i}$ affects the axial electric field. For that, we first introduce a vector field $\boldsymbol{v}$ (V), defined as
@@
Then, we define a relationship between the vector field $\boldsymbol{v}$ (V) and the axial electric field $\boldsymbol{e}$ (V/m). The axial electric field $\boldsymbol{e}$ describes how the voltage changes along one strand. Due to the periodicity of the problem, the integrated change of any quantity over a distance $\ell_{\text{s}}$ along a strand is equal to the change of that quantity between adjacent strands on the same cross-section. Hence, for crossing IS coupling currents, any net variation of the voltage $v_{y}$ along $\hat{\boldsymbol{e}}_{x}$ must be related to an axial electric field $\boldsymbol{e}$. This can be translated into the following equation:
@@
with $\bar{\boldsymbol{\sigma}}_{\text{IS}}$ (S/m3) the tensorial parameter (interpreted as an effective IS conductivity per cable area), defined by
@@ 4.1.2 Weak form of the DISCC model
The vector field $\boldsymbol{k}$ belongs to a function space $\mathcal{K}_{0}(\Omega_{\text{c}})$, that we define as
@@
The second term, a surface integral, is equal to zero, because $(\boldsymbol{j}\times\boldsymbol{n})\cdot\boldsymbol{k}^{\prime}=-(\boldsymbol{k}^{\prime}\times\boldsymbol{n})\cdot\boldsymbol{j}$, and $\boldsymbol{k}^{\prime}\times\boldsymbol{n}|_{\partial\Omega_{\text{c}}}=\boldsymbol{0}$ from the function space definition in Eq. (20). Using Ampère’s law $\boldsymbol{j}=\textrm{{curl}}\,\boldsymbol{h}$, the final weak form reads: from an initial solution at $t=0$, find $\boldsymbol{h}\in\mathcal{H}(\Omega)$ and $\boldsymbol{k}\in\mathcal{K}_{0}(\Omega_{\text{c}})$ such that, for $t>0$, $\forall\boldsymbol{h}^{\prime}\in\mathcal{H}_{0}(\Omega)$, and $\boldsymbol{k}^{\prime}\in\mathcal{K}_{0}(\Omega_{\text{c}})$,
@@ 5 Verification of the linear model
In this section, we evaluate the accuracy of the DISCC model in a linear setting, focusing on the IS coupling current dynamics only. We consider the reference cable model of Section 3.1 and the homogenized cable model of Section 4.1. The analysis is extended later in Section 6 by including in both the reference and the homogenized cable models the strand dynamics via the ROHM, ROHF, and current sharing models.
We first describe in Section 5.1 how to tune the values of the scaling parameters $\gamma_{\text{c}}$, $\gamma_{\text{a}}$, and $\lambda_{\text{a}}$, solely based on the calculated loss per cycle. In Section 5.2, we then discuss the accuracy of the homogenized model on both local fields and global quantities. Finally, we show in Section 5.3 how the results can be extrapolated for different cable properties.
Because the problem is linear, simulations can be performed in the frequency domain in terms of phasor fields. The phasor of the magnetic field is denoted by $\boldsymbol{\hat{h}}(\boldsymbol{x})$, with $\boldsymbol{x}$ the position vector, and is related to the physical magnetic field by $\boldsymbol{h}(\boldsymbol{x},t)=\Re\big(\boldsymbol{\hat{h}}(\boldsymbol{x})e^{i\omega t}\big)$, with $i=\sqrt{-1}$ and $\omega=2\pi f$, where $f$ is the frequency. The phasor for $\boldsymbol{k}$ is defined similarly and is denoted by $\boldsymbol{\hat{k}}$. Phasors are also introduced for currents and voltages in Eqs. (3.1) and (23), and time derivatives are replaced with $i\omega$.
For the reference model, the IS coupling current loss per cycle and per unit length $Q_{\text{IS}}$ (J/m) is directly computed as the sum of the Joule loss in all the lumped resistors of the electrical circuit of Fig. 4(c), that is,
$Q_{\text{IS}}=\frac{1}{2f\,\ell_{\text{s}}}\sum_{r\,\in\,\text{resistors}}{\hat{V}}_{r}^{\star}\hat{I}_{r},$ (24)
where ${\hat{I}}_{r}$ and ${\hat{V}}_{r}=R_{r}{\hat{I}}_{r}$ are the current and voltage phasors associated with resistor $r$ of resistance $R_{r}$, and where ${\hat{V}}^{\star}_{r}$ denotes the complex conjugate of $\hat{V}_{r}$.
@@
$Q_{\text{IS}}=\frac{1}{2f}\int_{\Omega_{\text{c}}}\boldsymbol{\hat{k}}^{\star}\cdot(\bar{\boldsymbol{\sigma}}_{\text{IS}}\,\boldsymbol{\hat{k}})\ \text{d}\Omega_{\text{c}},$ (25)
where $\boldsymbol{\hat{k}}^{\star}$ denotes the complex conjugate of $\boldsymbol{\hat{k}}$.
@@ 5.1 Parameter tuning: $\gamma_{\text{c}}$, $\gamma_{\text{a}}$, and $\lambda_{\text{a}}$
For illustration, we consider the flat cable with the properties given in Table 1. It has no keystone angle and the IS contact resistivities are assumed to be uniform over the cable. A uniform transverse magnetic field, described by a phasor $\boldsymbol{\hat{h}}_{\text{app}}=h_{\text{app}}\boldsymbol{\hat{\theta}}$, is applied in a direction $\boldsymbol{\hat{\theta}}$ that makes an angle $\theta$ with $\hat{\boldsymbol{e}}_{x}$. We fix $\mu_{0}h_{\text{app}}=10$ mT (the equations being linear, the actual amplitude does not matter). We cover frequencies ranging from $10$ mHz to $10$ kHz.
@@
For the main crossing parameter $\gamma_{\text{c}}$, we consider a field directed along $\hat{\boldsymbol{e}}_{y}$. Varying the value of $\gamma_{\text{c}}$ results in a frequency shift of the loss curve. The value of $\gamma_{\text{c}}$ can be adjusted in order to ensure that the homogenized loss curve best fits the reference loss curve. For the considered cable, the value $\gamma_{\text{c}}=0.43$ leads to an excellent match, at least for frequencies below $\approx 500$ Hz. The situation at higher frequencies is discussed later in Section 5.2. The relative difference between peak loss values is smaller than $0.5\%$. The match between the loss amplitudes and between the shape of the loss curves is a direct result of the DISCC model which requires no further tuning.
@@ 5.2 Result analysis
Using the scaling parameter values $\gamma_{\text{c}}=0.43$, $\gamma_{\text{a}}=0.53$, and $\lambda_{\text{a}}=0.006$, obtained in the previous section, we now analyze the accuracy of the DISCC model in detail.
5.2.1 Applied field along $\hat{\boldsymbol{e}}_{y}$
We first consider an external applied field $\boldsymbol{\hat{h}}_{\text{app}}=h_{\text{max}}\hat{\boldsymbol{e}}_{y}$ ($\theta=90\degree$). The complex field phasors are decomposed in their in-phase components, in phase with the applied field, i.e., they represent the solution when the applied field is maximum, and their quadrature components, which represent the solution a quarter of period later, i.e., when the applied field is zero.
@@ 5.2.1 Applied field along $\hat{\boldsymbol{e}}_{y}$
The current density distribution and magnetic flux density field lines computed by both the reference and homogenized models are illustrated in Fig. 11 for four frequencies. The $\boldsymbol{k}$ vector field distribution calculated by the homogenized model is also represented in Fig. 11. Magnetic flux density distributions of both models are compared in Fig. 12.
The agreement between both models is very good. The homogenized model correctly reproduces the field distortion induced by the IS coupling currents: as the frequency increases, the current density is progressively concentrated at the edges of the cable. Figure 12(b) highlights how the homogenized model produces averaged field distributions compared to the reference model inside the cable. Results of the reference model show slight inter-strand oscillations whereas results from the homogenized model are smoother.
The agreement between the homogenized and reference models worsens at higher frequencies, at which the current density distribution is concentrated in the strands at the edges of the cable (see Fig. 11 for $f=10$ kHz). Since the reference model uses strand-wise uniform current density while the homogenized model has no notion of discrete strands, a discrepancy is expected between both models at high frequencies. As will be shown in Section 6, strand loss contributions typically dominate at these frequencies (mainly eddy currents, see Fig. 21), such that the impact of this difference between the models on the total loss is small.

### v2

@@ 3.2 Nonlinear model with strand internal dynamics
The relation between strand current and voltage consists of two parts. The first part accounts for nonlinear inductive effects and is equal to the time derivative of a hysteretic internal magnetic flux $\Phi$ stored in the strand. The relationship between the strand current $I_{i}$ and the flux $\Phi$ is described by the ROHF model [20]. It is similar to the ROHM model, but it defines a scalar relationship and does not contain a branch for IF coupling currents, as these can be neglected for pure transport current excitations. The ROHF model contributes to the hysteresis and eddy loss. We refer to A.2 and [20] for details about the ROHF model and its implementation.
The second part of the current to voltage relationship reproduces resistive effects due to current sharing [38] between saturating superconducting filaments, whose resistivity is described by the power law [45], and the normal conducting strand. The current sharing solution [46] defines a nonlinear resistive contribution $E(I_{i})$ to the strand voltage, for a given current $I_{i}$. It is associated with the ohmic loss contribution. See A.3 for an example.
Introducing the ROHM and ROHF models, as well as the current sharing model, the formulation in Eq. (3.1) can be generalized as
@@
This problem is nonlinear and cannot be solved in the frequency domain anymore. An implicit Euler scheme is used for time-stepping, and a Newton-Raphson technique is implemented as an iterative scheme. The Jacobian tensor for the ROHM model is derived analytically, as in [19], and the derivatives $\text{d}\Phi/\text{d}I$ and $\text{d}E/\text{d}I$ can be easily derived analytically as well. The convergence criterion for the Newton-Raphson scheme is a condition on the relative change of the instantaneous power computed for two successive iterations.
@@ 4 Homogenized cable model
In this section, we introduce the homogenized cable model, see Figs. 1(f) and 6. The cable is no longer modelled with individual strands and circuit connections for IS coupling currents, but rather with a homogeneous material described with differential equations that reproduce the IS coupling current dynamics in a distributed manner. We refer to this model as the Distributed Inter-Strand Coupling Current (DISCC) model. A key assumption of this model is that strands are numerous enough for a description in terms of distributed field quantities to be sufficiently accurate.
Figure 6: FE meshes for reference and homogenized cable models. The homogenized cable domain $\Omega_{\text{c}}$ does not include the artificial coating domain $\Omega_{\text{coat}}$ of the reference model. The homogenized cable domain is the smallest rectangle (or trapezoid, see B) that contains all the bare strands, and the gaps between them, as represented in red in both subfigures.
In the homogenized cable model, equations are written in terms of homogenized fields, which are spatially averaged versions of the fields in the reference cable model. For simplicity, we keep the same notations for the fields as before, e.g., in this section, $\boldsymbol{h}$ and $\boldsymbol{j}$ refer to the homogenized magnetic field and current density.
Because contacts between strands are treated in a distributed way, we introduce contact resistivities $r_{\text{a}}$ and $r_{\text{c}}$ ($\Omega\,$m2), defined as the contact resistances multiplied by characteristic surfaces:
@@
We first describe the DISCC model in a linear setting in Section 4.1, focussing only on the IS coupling currents. We then generalize the model in Section 4.2 by including hysteresis, eddy, and IF coupling current effects via distributed versions of the ROHM, ROHF, and current sharing models. The accuracy and computational performance of the homogenized model is discussed in detail in Sections 5 and 6.
@@ 4.1 Linear model for distributed IS coupling currents
We start in a linear setting in which strands are considered to be non-magnetic and non-resistive. The IS coupling currents constitute the only loss contribution in this case.
@@ 4.1.1 Strong form of the DISCC model
For simplicity, we consider flat cables with a rectangular cross-section. We also assume uniform adjacent $r_{\text{a}}$ and crossing $r_{\text{c}}$ contact resistivities in the cable. These simplifications are however not necessary for the model, we discuss how to handle keystone angle effects [48] and non-uniform resistivities in B.
In absence of IS coupling currents, every strand carries the same net transport current (we assume that the current is uniformly shared among the strands at the termination joints), such that the homogenized current density $\boldsymbol{j}$ must be uniform.
@@
Figure 7: Two excitation modes for the IS coupling currents, with external field ramp $\dot{\boldsymbol{h}}_{\text{app}}$ applied along (a) $\hat{\boldsymbol{e}}_{y}$, or (b) $\hat{\boldsymbol{e}}_{x}$. This illustrates that, in first approximation, (a) a variation of the strand currents along $\hat{\boldsymbol{e}}_{x}$ is associated with crossing IS coupling currents along $\hat{\boldsymbol{e}}_{y}$, and that (b) a difference of the strand currents between both layers (variation along $\hat{\boldsymbol{e}}_{y}$) is associated with adjacent IS coupling currents along $\hat{\boldsymbol{e}}_{x}$ (strands in the top schematics are represented without adjacent contacts for better visualization). This justifies the use of the curl operator on the axial current density to describe IS coupling currents.
It is important to mention that Eq. (7) is only an approximation. Indeed, non-zero adjacent IS coupling currents (along $\hat{\boldsymbol{e}}_{x}$) also take place for magnetic field applied along $\hat{\boldsymbol{e}}_{y}$, as can be seen in Fig. 5 and as discussed in [49, 32]. We however keep Eq. (7) as an assumption, and we account for these adjacent IS currents later in this section (via the mixing parameter $\lambda_{\text{a}}$ in Eq. (15)).
Moreover, Eq. (7) has to be seen as a conceptual relationship, which encodes the geometrical periodicity of the cable in a distributed manner. The curl is not the expected operator (exterior derivative) for the current density (2-form) [50, 51] but for now, we focus on presenting intermediate steps leading to the weak form of the problem. In the final weak form, $\textrm{{curl}}\,\boldsymbol{j}$ does not appear anymore.
To close the mathematical problem, we must now express how the IS coupling current density field $\boldsymbol{i}$ affects the axial electric field. For that, we first introduce a vector field $\boldsymbol{v}$ (V), defined as
@@
Then, we need a relationship between the vector field $\boldsymbol{v}$ (V) and the axial electric field $\boldsymbol{e}$ (V/m). The axial electric field $\boldsymbol{e}$ describes how the voltage changes along one strand. Due to the periodicity of the problem, the integrated change of any quantity over a distance $\ell_{\text{s}}$ along a strand is equal to the change of that quantity between adjacent strands on the same cross-section. Hence, for crossing IS coupling currents, any net variation of the voltage $v_{y}$ along $\hat{\boldsymbol{e}}_{x}$ must be related to an axial electric field $\boldsymbol{e}$. This can be translated into the following equation:
@@
with $\bar{\boldsymbol{\sigma}}_{\text{IS}}$ (S/m3) a tensorial parameter (interpreted as an effective IS conductivity per cable area), defined by
@@ 4.1.2 Weak form of the DISCC model
The vector field $\boldsymbol{k}$ is purely transverse and belongs to a function space $\mathcal{K}_{0}(\Omega_{\text{c}})$, that we define as
@@
The second term, a surface integral, is equal to zero because $(\boldsymbol{j}\times\boldsymbol{n})\cdot\boldsymbol{k}^{\prime}=-(\boldsymbol{k}^{\prime}\times\boldsymbol{n})\cdot\boldsymbol{j}$, and $\boldsymbol{k}^{\prime}\times\boldsymbol{n}|_{\partial\Omega_{\text{c}}}=\boldsymbol{0}$ from the function space definition in Eq. (20). Using Ampère’s law $\boldsymbol{j}=\textrm{{curl}}\,\boldsymbol{h}$, the final weak form reads: from an initial solution at $t=0$, find $\boldsymbol{h}\in\mathcal{H}(\Omega)$ and $\boldsymbol{k}\in\mathcal{K}_{0}(\Omega_{\text{c}})$ such that, for $t>0$, $\forall\boldsymbol{h}^{\prime}\in\mathcal{H}_{0}(\Omega)$, and $\boldsymbol{k}^{\prime}\in\mathcal{K}_{0}(\Omega_{\text{c}})$,
@@ 5 Verification of the linear model
In this section, we evaluate the accuracy of the DISCC model in a linear setting, focusing on the IS coupling current dynamics only. We consider the reference cable model of Section 3.1 and the homogenized cable model of Section 4.1. The analysis is extended later in Section 6 by including in both the reference and the homogenized cable models the internal strand dynamics via the ROHM, ROHF, and current sharing models.
We first describe in Section 5.1 how to tune the values of the scaling parameters $\gamma_{\text{c}}$, $\gamma_{\text{a}}$, and $\lambda_{\text{a}}$, solely based on the calculated loss per cycle. In Section 5.2, we then discuss the physical results and the accuracy of the homogenized model in terms of local fields and global quantities. Finally, we show in Section 5.3 how the results can be extrapolated for different cable properties.
Because the problem is linear, simulations can be performed in the frequency domain using phasor fields. The phasor of the magnetic field is denoted by $\mathbf{h}(\boldsymbol{x})$, with $\boldsymbol{x}$ the position vector, and is related to the physical magnetic field by $\boldsymbol{h}(\boldsymbol{x},t)=\Re\big(\mathbf{h}(\boldsymbol{x})e^{i\omega t}\big)$, with $i=\sqrt{-1}$ and $\omega=2\pi f$, where $f$ is the frequency. The phasor for $\boldsymbol{k}$ is defined similarly and is denoted by $\mathbf{k}$. Phasors are also introduced for currents and voltages in Eqs. (3.1) and (23), and time derivatives are replaced with $i\omega$.
For the reference model, the IS coupling current loss per cycle and per unit length $Q_{\text{IS}}$ (J/m) is directly computed via the sum of the Joule loss in all the lumped resistors of the electrical circuit of Fig. 4(c), that is,
$Q_{\text{IS}}=\frac{1}{2f\,\ell_{\text{s}}}\sum_{r\,\in\,\text{resistors}}\mathrm{V}_{r}^{\star}\mathrm{I}_{r},$ (24)
where $\mathrm{I}_{r}$ and $\mathrm{V}_{r}=R_{r}\mathrm{I}_{r}$ are the current and voltage phasors associated with resistor $r$ of resistance $R_{r}$, and where $\mathrm{V}^{\star}_{r}$ denotes the complex conjugate of $\mathrm{V}_{r}$.
@@
$Q_{\text{IS}}=\frac{1}{2f}\int_{\Omega_{\text{c}}}\mathbf{k}^{\star}\cdot(\bar{\boldsymbol{\sigma}}_{\text{IS}}\,\mathbf{k})\ \text{d}\Omega_{\text{c}},$ (25)
where $\mathbf{k}^{\star}$ denotes the complex conjugate of $\mathbf{k}$.
@@ 5.1 Parameter tuning: $\gamma_{\text{c}}$, $\gamma_{\text{a}}$, and $\lambda_{\text{a}}$
For illustration, we consider the flat cable with the properties given in Table 1. It has no keystone angle and the IS contact resistivities are assumed to be uniform over the cable. A uniform transverse magnetic field, described by a phasor $\mathbf{h}_{\text{app}}=h_{\text{app}}\boldsymbol{\hat{\theta}}$, is applied in a direction $\boldsymbol{\hat{\theta}}$ that makes an angle $\theta$ with $\hat{\boldsymbol{e}}_{x}$. We fix $\mu_{0}h_{\text{app}}=10$ mT. We cover frequencies ranging from $10$ mHz to $10$ kHz.
@@
For the main crossing parameter $\gamma_{\text{c}}$, we consider a field directed along $\hat{\boldsymbol{e}}_{y}$. Varying the value of $\gamma_{\text{c}}$ results in a frequency shift of the loss curve. The value of $\gamma_{\text{c}}$ can be adjusted in order to ensure that the homogenized loss curve best fits the reference loss curve. For the considered cable, the value $\gamma_{\text{c}}=0.43$ leads to an excellent match, at least for frequencies below $\approx 500$ Hz. The situation at higher frequencies is discussed later in Section 5.2. The relative difference between peak loss values is smaller than $0.5\%$. The match between the loss amplitudes and between the shape of the loss curves is a direct result of the DISCC model which requires no tuning.
@@ 5.2 Result analysis
Using the scaling parameter values $\gamma_{\text{c}}=0.43$, $\gamma_{\text{a}}=0.53$, and $\lambda_{\text{a}}=0.006$, obtained in the previous section, we now analyze the accuracy of the DISCC model in detail and interpret the results.
@@
5.2.1 Applied field along $\hat{\boldsymbol{e}}_{y}$
We first consider an external applied field $\mathbf{h}_{\text{app}}=h_{\text{app}}\hat{\boldsymbol{e}}_{y}$ ($\theta=90\degree$). The complex field phasors are decomposed in their in-phase components, in phase with the applied field, i.e., they represent the solution when the applied field is maximum, and their quadrature components, which represent the solution a quarter of period later, i.e., when the applied field is zero.
The current density distribution and magnetic flux density field lines computed by the reference and homogenized models are illustrated in Fig. 11 for four frequencies. The $\boldsymbol{k}$ vector field distribution calculated by the homogenized model is also represented in Fig. 11. Magnetic flux density distributions of both models are compared in Fig. 12. In particular, Fig. 12(b) highlights how the homogenized model produces averaged field distributions compared to the reference model inside the cable. Results of the reference model show slight inter-strand oscillations whereas results from the homogenized model are smoother.
The agreement between both models is very good. The homogenized model correctly reproduces the progressive field distortion induced by the IS coupling currents. At low frequencies ($f\lesssim 0.2$ Hz), the amplitude of the induced IS coupling currents is too small to significantly distort the applied field and, as a result, the IS coupling current amplitude is almost proportional to $f$, such that the instantaneous loss curve scales with $f^{2}$, and the loss per cycle with $f$.
When the frequency increases further, the field generated by the IS coupling currents gets comparable with the applied field and the loss per cycle reaches a peak, after which it starts to decrease. This is caused by the current density being progressively concentrated towards the edges of the cable, as illustrated in Fig. 11. The exact evolution of the loss per cycle and field lines with frequency is a result of the cable shape and of the equations governing the IS coupling current dynamics. Both models produce very similar loss evolution up to $f\lesssim 500$ Hz, as shown in Fig. 8.
The agreement between the models worsens at higher frequencies for which the current density distribution is concentrated in the strands at the edges of the cable (see Fig. 11 for $f=10$ kHz). Since the reference model uses strand-wise uniform current density, it is incapable or representing current density concentration smaller than the size of a strand. On the other hand, the homogenized model has no notion of discrete strands. A discrepancy is therefore expected between both models at high frequencies. However, as will be shown in Section 6, other loss contributions typically dominate at these frequencies, such as eddy currents and ohmic effects, such that the impact of this difference between the models on the total loss is small and acceptable.
The influence of the spatial discretization on $Q_{\text{IS}}$ is illustrated in Fig. 13. From mesh $2$ to mesh $5$, the relative difference between calculated peak losses is less than $1.8\%$. This indicates that a relatively coarse mesh can be chosen in practice, e.g., mesh 2 or 3. The difference between the results increases with frequency, because the field distributions become increasingly non-uniform and concentrated at the edges of the cable when the frequency increases (see Fig. 11), such that the effect of the spatial discretization is visible.